in a triangle ABC, side a is twice as long as side b. If the hypotenuse is 30 m longe,solve
Answer by checkley77(12844)

a^2+(2a)^2=30^2
A^2+4A^2=900
5A^2=900
A^2=900/5
A^2=180
A=SQRT180
A=13.4164M. ANS. FOR THE SHORT SIDE.
2*13.4164=26.8328M. ANS. FOR THE LONGEST SIDE.
PROOF:
13.4164^2+26.8328^2=900
180+720=900
900=900